OpenStudy (anonymous):

If 5 sin theta = 3. Find sec theta+tan theta/sec theta - tan theta

OpenStudy (anonymous):

\[\sin(\theta)=\frac{3}{5}\] so you have a right triangle with opposite side 3, adjacent side 4 and hypotenuse 5

OpenStudy (anonymous):

draw the triangle and you can find all other trig funcitons

OpenStudy (anonymous):

thanx! :D

OpenStudy (anonymous):

if you get stuck on any let me know

OpenStudy (anonymous):

can you solve and show if you dont mind?

OpenStudy (anonymous):

\[\sin \theta = 3/5 =opposite/hypotenuse\]

OpenStudy (anonymous):

use pythagoras theorem to find the third side

OpenStudy (anonymous):

ok so your hypotenuse is 5, opposite is 3 adjacent is 4 here are all of them right from the triangle

OpenStudy (anonymous):

\[\sin(\theta)=\frac{3}{5}\] \[\cos(\theta)=\frac{4}{5}\] \[\tan(\theta)=\frac{3}{4}\]

OpenStudy (anonymous):

now for the reciprocals \[\csc(\theta)=\frac{5}{3}\] \[\sec(\theta)=\frac{5}{4}\] \[\cot(\theta)=\frac{4}{3}\]

OpenStudy (anonymous):

notice how odd this is because you are asked to evaluate a function without actually knowing what the input is. you rely on the ratio of the triangle to get it, once you have memorized that whole "opposite over hypotenuse" business